# GEBench: Benchmarking Image Generation Models as GUI Environments

### v1

@@ 1 Introduction
Recent advancements in image generation models hurst2024gpt; comanici2025gemini; team2023gemini; seedream2025seedream; flux-2-2025; wan2025 enable the prediction of future Graphical User Interface (GUI) states based on specific user instructions and current visual contexts. This capability positions image generation models as powerful GUI Environments zhang2025large; yan2025gui; xie2025mirage; garg2025controllable; luo2025vimo; wei2023boosting, capable of simulating dynamic interaction sequences to facilitate the scalable training of autonomous agents. Distinguished from conventional simulators cobbe2020leveraging; xie2024osworld; bonatti2024windows tethered to physical hardware or fixed software stacks zhang2025large, these generative models offer a flexible, low-cost alternative for creating diverse interaction trajectories across countless applications zhao2021guigan; liu2025ui.
@@
However, the potential of image generation models as reliable GUI environments remains largely unverified, as traditional visual benchmarks ghosh2023geneval; hu2024ella; niu2025wise; zhao2025envisioning; huang2023t2i; huang2024vbench; sun2025t2v; zhuang2025vistorybench prioritize general-domain visual fidelity (for images) and continuous state transitions (for videos), leaving a critical gap in evaluating the functional logic and state-transition consistency inherent to GUI interactions xie2025gui; yan2025gui. As shown in Figure 1, when acting as GUI environments zhang2025large, generation models must seamlessly navigate discrete, action-triggered interface jumps. Such transitions necessitate precise coordinate grounding zhao2021guigan; cheng2024seeclick, icon recognition liu2025ui; xie2025gui, and high-fidelity text rendering chen2024textdiffuser, compelling the models to maintain logical continuity even when visual elements do not persist li2025mobileworldbench. Such demands strain existing architectures and call for a new evaluation approach to verify if generated GUIs respond felicitously to user instructions.
To bridge this gap, we present GEBench (Benchmarking image generation models as GUI Environments), a benchmark designed to evaluate how effectively image generation models can serve as GUI environments. GEBench comprises 700 high-quality samples, where each entry aligns a sequence of GUI images with corresponding user instructions. These samples span five distinct tasks, allowing for a multifaceted assessment of the model’s ability. To provide a concrete measure of generative quality, we propose GE-Score, a multi-dimensional metric derived from Vision Language Model (VLM)-guided google2025gemini3; hurst2024gpt; bai2025qwen3vltechnicalreport evaluations across five specialized rubrics. GE-Score systematically validates intent fulfillment and interaction logic while verifying UI content consistency and structural integrity. By ensuring high visual fidelity and logical coherence, GE-Score confirms the practical utility of these synthetic environments.
Our systematic evaluation of state-of-the-art image generation models openai2025gptimage; team2023gemini; seedream2025seedream; seedream2025seed; flux-2-2025; wan2025; wu2025qwen; deng2025emerging; li2025uniworld; team2025longcat identifies promising pathways for their evolution into reliable GUI environments. While current architectures demonstrate robust proficiency in executing localized, single-step state transitions, they offer significant opportunities for advancement in long-term interaction consistency and precise spatial grounding. In particular, deficiencies in icon interpretation, Chinese text rendering, and grounding point localization lead to layout drift and logical inconsistencies. These observations delineate critical bottlenecks and outline clear directions for future research toward high-fidelity, temporally coherent generative GUI systems.
@@ 2.1 Automated GUIs Generation
The evolution of GUIs generation reflects a significant paradigm shift from heuristic-based structural mapping to data-driven synthesis powered by Multimodal Large Language Models (MLLMs) chen2018ui; sandhaus2011employing; yang2016automatic; li2020layoutgan; zhao2021guigan; mozaffari2022ganspiration; sobolevsky2023guilget; zhang2025scaling; kolthoff2025guide; kolthoff2024zero. Early methodologies relied on traditional rule-based algorithms to perform layout reconstruction sandhaus2011employing; huang2016automaticly, yet these approaches frequently failed to capture the semantic depth of complex hierarchies. Subsequent frameworks simplified this process using model-based approaches to translate visual features directly into code sequences chen2018ui. Contemporary research leverages Transformer-based architectures to bridge the gap between visual design abstractions and executable source code sobolevsky2023guilget; kolthoff2025guide. Furthermore, the rapid advancement of generative AI suggests that direct utilization of image generation models for GUI synthesis is becoming increasingly viable li2020layoutgan; zhao2021guigan; mozaffari2022ganspiration; zhang2025latent. These models offer the potential to produce high-fidelity GUIs directly from user instructions.
@@ 2.2 Advanced Image Generation Models
Recent progress in image generation exhibits a rapid evolution from text-to-image synthesis to sophisticated reference-based frameworks hurst2024gpt; comanici2025gemini; team2023gemini; seedream2025seedream; wan2025; deng2025emerging; liu2025step1x; li2025uniworld; team2025nextstep; team2025longcat. Ongoing advancements in text-to-image synthesis have empowered models to produce aesthetically superior visuals with precise semantic alignment to the provided instructions ho2020denoising; chen2020generative; flux-2-2025; fan2024fluid; han2025infinity; ramesh2022hierarchical; lin2025perceiveanythingrecognizeexplain. Building on these foundations, reference-based techniques integrate visual priors with textual prompts to enhance generative control team2023gemini; seedream2025seedream; wan2025; team2025longcat; an2025unictokens. These methods incorporate style or structural references to ensure spatial precision and identity consistency deng2025emerging; liu2025step1x. These advances enable image generation models to function as interactive GUI environments.
@@ 2.3 Sequential Generation Benchmarks
Standard image generation benchmarks focus on visual fidelity and text-alignment for a single image, using metrics like FID and CLIP score to measure visual quality heusel2017gans; ghosh2023geneval; huang2023t2i; clip_metrics2021. While these metrics provide a robust baseline for aesthetic realism heusel2017gans, they set the stage for incorporating logical coherence to achieve a more holistic evaluation. Recent efforts in benchmarking sequential image generation explore temporal consistency and reasoning niu2025wise; hu2024ella; zhao2025envisioning; huang2023t2i; zhuang2025vistorybench; zhang2026vibe-benchmark; guo2025video, yet these frameworks typically target natural scenes with continuous movement, simple spatial relationships or characters identity liu2025step1x. GUI environments differ fundamentally because they involve discrete state jumps where a single action replaces the entire visual layout zhang2025large; yan2025gui. Furthermore, the stringent text-rendering requirements of GUIs push current generative architectures to their limits chen2024textdiffuser. This necessitates a new evaluation approach to bridge the significant gap in assessing whether image generation models can maintain the strict semantic and structural integrity required for multi-step GUI trajectories generation.
@@ 3 GEBench
A GUI environment functions as an interactive medium that translates user instructions or agent actions into corresponding visual feedback through software logic zhang2025large. This mechanism allows a system to simulate the evolution of a digital interface in response to user interventions. The highly structured and rule-based composition of GUIs, governed by precise functional logic, distinguishes these environments from the patterns of natural scenes.
@@ 3.3 Data Construction Pipeline
Nano Banana pro team2023gemini 84.50 68.65 65.75 64.35 64.83 69.62 84.32 69.51 46.33 47.20 58.64 61.20
Nano Banana team2023gemini 64.36 34.16 64.82 65.89 54.48 56.74 64.80 50.75 48.88 47.12 49.04 52.12
GPT-image-1.5 openai2025gptimage 83.79 56.97 60.11 55.65 53.33 63.22 80.80 58.87 63.68 58.93 49.23 63.16
GPT-image-1.0 openai2025gptimage 64.72 49.20 57.31 59.04 31.68 52.39 60.92 64.33 58.94 56.16 37.84 55.64
Seedream 4.5 seedream2025seed 63.64 53.11 56.48 53.44 52.90 55.91 49.49 45.30 53.81 51.80 49.63 50.01
Seedream 4.0 seedream2025seedream 62.04 48.64 49.28 50.93 53.53 52.88 53.28 37.57 47.92 49.36 44.17 46.46
Wan 2.6 wan2025 64.20 50.11 52.72 50.40 59.58 55.40 60.17 44.36 49.55 44.80 53.36 50.45
Flux-2-pro flux-2-2025 68.83 55.07 58.13 55.41 50.24 57.54 61.00 52.17 49.92 47.16 45.67 51.18
Bagel deng2025emerging 34.84 13.45 27.36 33.52 35.10 28.85 32.91 8.61 26.08 35.12 37.30 28.00
UniWorld-V2 li2025uniworld 55.33 24.95 32.03 21.39 49.60 36.66 42.68 14.14 30.08 26.83 47.04 32.15
Qwen-Image-Edit wu2025qwen 41.12 26.79 23.78 26.10 50.80 33.72 40.12 18.61 25.80 25.95 54.55 33.01
Longcat-Image team2025longcat 48.76 12.75 30.03 30.00 51.02 34.51 36.69 8.44 37.30 36.83 47.12 33.28
@@ 4.1 Evaluation Setup
The commercial model group includes Nano Banana pro team2023gemini, Nano Banana team2023gemini, GPT-image-1.5 openai2025gptimage, GPT-image-1 openai2025gptimage, Seedream 4.5 seedream2025seed, Seedream 4.0 seedream2025seedream, Wan2.6 wan2025, Flux-2-pro flux-2-2025.
The open-source model group includes Bagel deng2025emerging, UniWorld-V2 li2025uniworld, Qwen-Image-Edit wu2025qwen, LongCat-Image team2025longcat.
@@
To ensure the objectivity and robustness of GEBench, we deploy 3 state-of-the-art VLMs as independent cross-evaluators: 2 commercial models Gemini-3-Flash-Native google2025gemini3, GPT-4o hurst2024gpt and 1 open-source model Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport. By utilizing these evaluators, we mitigate potential bias inherent in a single judge model. To ensure fair and reproducible comparisons, we use official default configurations for evaluated models and perform evaluation three times for each generated GUIs trajectory.
@@ 4.2 Evaluation Results
Experimental results in Table 1 show that Nano Banana Pro team2023gemini delivers the most robust performance, particularly on Chinese subset with a top-ranking GE-Score of 69.62. GPT-image-1.5 openai2025gptimage follows closely, excelling on English subset and securing the first position with a score of 63.16. The radar chart, as shown in Figure 4, further illustrates that commercial models, led by Nano Banana Pro team2023gemini, exhibit a balanced and “full" pentagonal profile. In contrast, open-source models (e.g., UniWorld-V2 li2025uniworld, Bagel deng2025emerging) show performance curves that significantly shrink inward, revealing a substantial gap in handling complex tasks.
@@
The evaluation results in Table 1 highlights a critical bottleneck: while most models excel in Single-step transitions, with both Nano Banana Pro team2023gemini and GPT-image-1.5 openai2025gptimage exceeding 80 points, their scores plummet in Multi-step Planning scenarios, generally dropping below 60 or even 10 points. The radar chart, as shown in Figure 4, clearly identifies the Multi-step axis as a general weakness across nearly all models.
@@
Cross-metric analysis reveals a general deficit in Grounding-based Generation tasks. As shown in Figure 5, the GOAL (Goal Achievement) score is remarkably low: even top-performing Nano Banana Pro team2023gemini achieves only 23.9%, while most other models (e.g., Qwen-Image-Edit wu2025qwen) fall below 10%. The pronounced “dent" on the Grounding axis of the radar chart, as in Figure 4, further validates this bottleneck: The dismal GOAL scores indicate a logical disconnect; models identify what to generate but cannot translate this into where to place it on a precise $[0,1000]$ coordinate grid. This suggests models lack a fundamental understanding of the mapping between abstract coordinates and image pixels.
@@
Figure 6: Pearson correlation analysis between human expert scores and VLM-based evaluation. Results for Nano Banana Pro team2023gemini and GPT-Image-1 openai2025gptimage demonstrate a strong alignment between the VLM-as-a-Judge framework and human judgment across different models.
@@ 4.3 Validity of VLM-as-a-Judge
To validate the reliability of using VLM as evaluators, we analyze the correlation between VLM-based assessments and human expert judgments. Specifically, we randomly sample results from two representative models, Nano Banana Pro team2023gemini and GPT-Image-1 openai2025gptimage. For each model, we select 10 edited samples from each of the 10 tasks, resulting in 100 evaluated samples per model. Four human experts independently assess all selected samples using the same evaluation criteria and metrics as the VLM-based judges. Human scores are obtained by averaging all scores across four experts. We then compute the Pearson Correlation Coefficient between the human-annotated scores and the scores produced by the VLM-based judges.
As shown in Figure 6, our VLM-based evaluations exhibit strong correlation with human judgments. The overall Pearson correlation coefficient across all samples reaches r = 0.9892. When analyzed by model, the correlation remains consistently high, with r = 0.9926 for Nano Banana Pro team2023gemini and r = 0.9833 for GPT-Image-1 openai2025gptimage. These results indicate a high level of agreement between the VLM evaluator and human experts across different tasks and models. Together, this analysis demonstrates that the proposed VLM-as-a-Judge framework provides reliable and human-aligned evaluations across diverse tasks and models.

### v2

@@ 1 Introduction
Recent advancements in image generation models (hurst2024gpt; comanici2025gemini; team2023gemini; seedream2025seedream; flux-2-2025; wan2025) enable the prediction of future Graphical User Interface (GUI) states based on specific user instructions and current visual contexts. This capability positions image generation models as powerful GUI Environments (zhang2025large; yan2025gui; xie2025mirage; garg2025controllable; luo2025vimo; wei2023boosting), capable of simulating dynamic interaction sequences to facilitate the scalable training of autonomous agents. Distinguished from conventional simulators (cobbe2020leveraging; xie2024osworld; bonatti2024windows) tethered to physical hardware or fixed software stacks (zhang2025large), these generative models offer a flexible, low-cost alternative for creating diverse interaction trajectories across countless applications (zhao2021guigan; liu2025ui).
@@
However, the potential of image generation models as reliable GUI environments remains largely unverified, as traditional visual benchmarks (ghosh2023geneval; hu2024ella; niu2025wise; zhao2025envisioning; huang2023t2i; huang2024vbench; sun2025t2v; zhuang2025vistorybench) prioritize general-domain visual fidelity (for images) and continuous state transitions (for videos), leaving a critical gap in evaluating the functional logic and state-transition consistency inherent to GUI interactions (xie2025gui; yan2025gui). As shown in Figure 1, when acting as GUI environments (zhang2025large), generation models must seamlessly navigate discrete, action-triggered interface jumps. Such transitions necessitate precise coordinate grounding (zhao2021guigan; cheng2024seeclick), icon recognition (liu2025ui; xie2025gui), and high-fidelity text rendering (chen2024textdiffuser), compelling the models to maintain logical continuity even when visual elements do not persist (li2025mobileworldbench). Such demands strain existing architectures and call for a new evaluation approach to verify if generated GUIs respond felicitously to user instructions.
To bridge this gap, we present GEBench (Benchmarking image generation models as GUI Environments), a benchmark designed to evaluate how effectively image generation models can serve as GUI environments. GEBench comprises 700 high-quality samples, where each entry aligns a sequence of GUI images with corresponding user instructions. These samples span five distinct tasks, allowing for a multifaceted assessment of the model’s ability. To provide a concrete measure of generative quality, we propose GE-Score, a multi-dimensional metric derived from Vision Language Model (VLM)-guided (google2025gemini3; hurst2024gpt; bai2025qwen3vltechnicalreport) evaluations across five specialized rubrics. GE-Score systematically validates intent fulfillment and interaction logic while verifying UI content consistency and structural integrity. By ensuring high visual fidelity and logical coherence, GE-Score confirms the practical utility of these synthetic environments.
Our systematic evaluation of state-of-the-art image generation models (openai2025gptimage; team2023gemini; seedream2025seedream; seedream2025seed; flux-2-2025; wan2025; wu2025qwen; deng2025emerging; li2025uniworld; team2025longcat) identifies promising pathways for their evolution into reliable GUI environments. While current architectures demonstrate robust proficiency in executing localized, single-step state transitions, they offer significant opportunities for advancement in long-term interaction consistency and precise spatial grounding. In particular, deficiencies in icon interpretation, Chinese text rendering, and grounding point localization lead to layout drift and logical inconsistencies. These observations delineate critical bottlenecks and outline clear directions for future research toward high-fidelity, temporally coherent generative GUI systems.
@@ 2.1 Automated GUIs Generation
The evolution of GUIs generation reflects a significant paradigm shift from heuristic-based structural mapping to data-driven synthesis powered by Multimodal Large Language Models (MLLMs) (chen2018ui; sandhaus2011employing; yang2016automatic; li2020layoutgan; zhao2021guigan; mozaffari2022ganspiration; sobolevsky2023guilget; zhang2025scaling; kolthoff2025guide; kolthoff2024zero). Early methodologies relied on traditional rule-based algorithms to perform layout reconstruction (sandhaus2011employing; huang2016automaticly), yet these approaches frequently failed to capture the semantic depth of complex hierarchies. Subsequent frameworks simplified this process using model-based approaches to translate visual features directly into code sequences (chen2018ui). Contemporary research leverages Transformer-based architectures to bridge the gap between visual design abstractions and executable source code (sobolevsky2023guilget; kolthoff2025guide). Furthermore, the rapid advancement of generative AI suggests that direct utilization of image generation models for GUI synthesis is becoming increasingly viable (li2020layoutgan; zhao2021guigan; mozaffari2022ganspiration; zhang2025latent). These models offer the potential to produce high-fidelity GUIs directly from user instructions.
@@ 2.2 Advanced Image Generation Models
Recent progress in image generation exhibits a rapid evolution from text-to-image synthesis to sophisticated reference-based frameworks (hurst2024gpt; comanici2025gemini; team2023gemini; seedream2025seedream; wan2025; deng2025emerging; liu2025step1x; li2025uniworld; team2025nextstep; team2025longcat). Ongoing advancements in text-to-image synthesis have empowered models to produce aesthetically superior visuals with precise semantic alignment to the provided instructions (ho2020denoising; chen2020generative; flux-2-2025; fan2024fluid; han2025infinity; ramesh2022hierarchical; lin2025perceiveanythingrecognizeexplain). Building on these foundations, reference-based techniques integrate visual priors with textual prompts to enhance generative control (team2023gemini; seedream2025seedream; wan2025; team2025longcat; an2025unictokens). These methods incorporate style or structural references to ensure spatial precision and identity consistency (deng2025emerging; liu2025step1x). These advances enable image generation models to function as interactive GUI environments.
@@ 2.3 Sequential Generation Benchmarks
Standard image generation benchmarks focus on visual fidelity and text-alignment for a single image, using metrics like FID and CLIP score to measure visual quality (heusel2017gans; ghosh2023geneval; huang2023t2i; clip_metrics2021). While these metrics provide a robust baseline for aesthetic realism (heusel2017gans), they set the stage for incorporating logical coherence to achieve a more holistic evaluation. Recent efforts in benchmarking sequential image generation explore temporal consistency and reasoning (niu2025wise; hu2024ella; zhao2025envisioning; huang2023t2i; zhuang2025vistorybench; zhang2026vibe-benchmark; guo2025video), yet these frameworks typically target natural scenes with continuous movement, simple spatial relationships or characters identity (liu2025step1x). GUI environments differ fundamentally because they involve discrete state jumps where a single action replaces the entire visual layout (zhang2025large; yan2025gui). Furthermore, the stringent text-rendering requirements of GUIs push current generative architectures to their limits (chen2024textdiffuser). This necessitates a new evaluation approach to bridge the significant gap in assessing whether image generation models can maintain the strict semantic and structural integrity required for multi-step GUI trajectories generation.
@@ 3 GEBench
A GUI environment functions as an interactive medium that translates user instructions or agent actions into corresponding visual feedback through software logic (zhang2025large). This mechanism allows a system to simulate the evolution of a digital interface in response to user interventions. The highly structured and rule-based composition of GUIs, governed by precise functional logic, distinguishes these environments from the patterns of natural scenes.
@@ 3.3 Data Construction Pipeline
Nano Banana pro (team2023gemini) 84.50 68.65 65.75 64.35 64.83 69.62 84.32 69.51 46.33 47.20 58.64 61.20
Nano Banana (team2023gemini) 64.36 34.16 64.82 65.89 54.48 56.74 64.80 50.75 48.88 47.12 49.04 52.12
GPT-image-1.5 (openai2025gptimage) 83.79 56.97 60.11 55.65 53.33 63.22 80.80 58.87 63.68 58.93 49.23 63.16
GPT-image-1.0 (openai2025gptimage) 64.72 49.20 57.31 59.04 31.68 52.39 60.92 64.33 58.94 56.16 37.84 55.64
Seedream 4.5 (seedream2025seed) 63.64 53.11 56.48 53.44 52.90 55.91 49.49 45.30 53.81 51.80 49.63 50.01
Seedream 4.0 (seedream2025seedream) 62.04 48.64 49.28 50.93 53.53 52.88 53.28 37.57 47.92 49.36 44.17 46.46
Wan 2.6 (wan2025) 64.20 50.11 52.72 50.40 59.58 55.40 60.17 44.36 49.55 44.80 53.36 50.45
Flux-2-pro (flux-2-2025) 68.83 55.07 58.13 55.41 50.24 57.54 61.00 52.17 49.92 47.16 45.67 51.18
Bagel (deng2025emerging) 34.84 13.45 27.36 33.52 35.10 28.85 32.91 8.61 26.08 35.12 37.30 28.00
UniWorld-V2 (li2025uniworld) 55.33 24.95 32.03 21.39 49.60 36.66 42.68 14.14 30.08 26.83 47.04 32.15
Qwen-Image-Edit (wu2025qwen) 41.12 26.79 23.78 26.10 50.80 33.72 40.12 18.61 25.80 25.95 54.55 33.01
Longcat-Image (team2025longcat) 48.76 12.75 30.03 30.00 51.02 34.51 36.69 8.44 37.30 36.83 47.12 33.28
@@ 4.1 Evaluation Setup
The commercial model group includes Nano Banana pro (team2023gemini), Nano Banana (team2023gemini), GPT-image-1.5 (openai2025gptimage), GPT-image-1 (openai2025gptimage), Seedream 4.5 (seedream2025seed), Seedream 4.0 (seedream2025seedream), Wan2.6 (wan2025), Flux-2-pro (flux-2-2025).
The open-source model group includes Bagel (deng2025emerging), UniWorld-V2 (li2025uniworld), Qwen-Image-Edit (wu2025qwen), LongCat-Image (team2025longcat).
@@
To ensure the objectivity and robustness of GEBench, we deploy 3 state-of-the-art VLMs as independent cross-evaluators: 2 commercial models Gemini-3-Flash-Native (google2025gemini3), GPT-4o (hurst2024gpt) and 1 open-source model Qwen3-vl-235b-a22b-thinking (bai2025qwen3vltechnicalreport). By utilizing these evaluators, we mitigate potential bias inherent in a single judge model. To ensure fair and reproducible comparisons, we use official default configurations for evaluated models and perform evaluation three times for each generated GUIs trajectory.
@@ 4.2 Evaluation Results
Experimental results in Table 1 show that Nano Banana Pro (team2023gemini) delivers the most robust performance, particularly on Chinese subset with a top-ranking GE-Score of 69.62. GPT-image-1.5 (openai2025gptimage) follows closely, excelling on English subset and securing the first position with a score of 63.16. The radar chart, as shown in Figure 4, further illustrates that commercial models, led by Nano Banana Pro (team2023gemini), exhibit a balanced and “full" pentagonal profile. In contrast, open-source models (e.g., UniWorld-V2 (li2025uniworld), Bagel (deng2025emerging)) show performance curves that significantly shrink inward, revealing a substantial gap in handling complex tasks.
@@
The evaluation results in Table 1 highlights a critical bottleneck: while most models excel in Single-step transitions, with both Nano Banana Pro (team2023gemini) and GPT-image-1.5 (openai2025gptimage) exceeding 80 points, their scores plummet in Multi-step Planning scenarios, generally dropping below 60 or even 10 points. The radar chart, as shown in Figure 4, clearly identifies the Multi-step axis as a general weakness across nearly all models.
@@
Cross-metric analysis reveals a general deficit in Grounding-based Generation tasks. As shown in Figure 5, the GOAL (Goal Achievement) score is remarkably low: even top-performing Nano Banana Pro (team2023gemini) achieves only 23.9%, while most other models (e.g., Qwen-Image-Edit (wu2025qwen)) fall below 10%. The pronounced “dent" on the Grounding axis of the radar chart, as in Figure 4, further validates this bottleneck: The dismal GOAL scores indicate a logical disconnect; models identify what to generate but cannot translate this into where to place it on a precise $[0,1000]$ coordinate grid. This suggests models lack a fundamental understanding of the mapping between abstract coordinates and image pixels.
@@
Figure 6: Pearson correlation analysis between human expert scores and VLM-based evaluation. Results for Nano Banana Pro (team2023gemini) and GPT-Image-1 (openai2025gptimage) demonstrate a strong alignment between the VLM-as-a-Judge framework and human judgment across different models.
@@ 4.3 Validity of VLM-as-a-Judge
To validate the reliability of using VLM as evaluators, we analyze the correlation between VLM-based assessments and human expert judgments. Specifically, we randomly sample results from two representative models, Nano Banana Pro (team2023gemini) and GPT-Image-1 (openai2025gptimage). For each model, we select 10 edited samples from each of the 10 tasks, resulting in 100 evaluated samples per model. Four human experts independently assess all selected samples using the same evaluation criteria and metrics as the VLM-based judges. Human scores are obtained by averaging all scores across four experts. We then compute the Pearson Correlation Coefficient between the human-annotated scores and the scores produced by the VLM-based judges.
As shown in Figure 6, our VLM-based evaluations exhibit strong correlation with human judgments. The overall Pearson correlation coefficient across all samples reaches r = 0.9892. When analyzed by model, the correlation remains consistently high, with r = 0.9926 for Nano Banana Pro (team2023gemini) and r = 0.9833 for GPT-Image-1 (openai2025gptimage). These results indicate a high level of agreement between the VLM evaluator and human experts across different tasks and models. Together, this analysis demonstrates that the proposed VLM-as-a-Judge framework provides reliable and human-aligned evaluations across diverse tasks and models.